OpenStudy (anonymous):

Simplify: (−12x^3y^4 + 6x^2y^2) ÷ 6x^2y^2

ganeshie8 (ganeshie8):

\(\large \frac{−12x^3y^4 + 6x^2y^2}{6x^2y^2}\)

ganeshie8 (ganeshie8):

start by factoring the GCF from numerator

ganeshie8 (ganeshie8):

and see if something cancels away wid the denominator

OpenStudy (anonymous):

I thought I could cancel out both 6x^2y^2

ganeshie8 (ganeshie8):

yes !!

ganeshie8 (ganeshie8):

\(\large \frac{−12x^3y^4 + 6x^2y^2}{6x^2y^2}\) factor out 6x^2y^2 in numerator \(\large \frac{6x^2y^2(−2xy^2 + 1)}{6x^2y^2}\)

ganeshie8 (ganeshie8):

now u can cancel out

ganeshie8 (ganeshie8):

that gives u \(\large -2xy^2 + 1\)

OpenStudy (anonymous):

So why did the 12x^2y^4 turn into -2xy^2 + 1 ?

OpenStudy (anonymous):

Anyway thank you :)

ganeshie8 (ganeshie8):

good question :)

ganeshie8 (ganeshie8):

\(\large \frac{−12x^3y^4 + 6x^2y^2}{6x^2y^2}\)

ganeshie8 (ganeshie8):

before cenceling, u need to take GCF of numerator

ganeshie8 (ganeshie8):

u cannot cancel the second term and denominator directly okay ?

ganeshie8 (ganeshie8):

GCF of numerator is \(6x^2y^2\)

ganeshie8 (ganeshie8):

when u pull it from numerator's first term, \(-12x^3y^4\), it leaves \(-2xy^2\)

ganeshie8 (ganeshie8):

\(\large -12x^3y^4 = 6x^2y^2(-2xy^2)\)

OpenStudy (anonymous):

ooooh I got it, I forgot about that.

OpenStudy (anonymous):

Thanks again.

ganeshie8 (ganeshie8):

good to hear :) u wlc :)